OpenStudy (anonymous):

The line that passes through (1, 5) and (-2, 3)

OpenStudy (anonymous):

first we have to find the slope and then have to use the points to find the y=mx + b equation.

OpenStudy (anonymous):

m (slope) = y2 - y1 / x2 - x1 so that is (3-5)/(-2-1) = -2/-3 = 2/3

OpenStudy (anonymous):

so now we have to use on of the points ( we will use 1,5) to write the equation in point slope form: y - y1 = m ( x-x1)

OpenStudy (anonymous):

that is y - 5 = (2/3)(x-1) and then put the problem in y=mx + b form and you have the equation for your line.